# PostEdit: Posterior Sampling for Efficient Zero-Shot Image Editing

### v1

@@ 3.2 Posterior Sampling for Image Editing
In this section, we provide a detailed illustration of the sampling process for high-quality image editing using the DDIM solver Luo et al. (2023a), as shown in Eq. 6. Unlike the ODE solver and measurement $\boldsymbol{y}$ used in the image reconstruction task described in Sec. 3.1, image editing requires a masked measurement derived from $\boldsymbol{z}_{0}$ and an ODE solver with higher accuracy to estimate $\boldsymbol{\hat{z}}_{0}$. The measurement $\boldsymbol{y}$ for image editing is defined as
@@
where $\boldsymbol{P}\in\{0,1\}^{n\times p}$ represents a masking matrix composed of elementary unit vectors. This measurement setup not only serves as a specialized configuration for image editing but also demonstrates its capacity to deliver high-quality image reconstruction results, even when $\boldsymbol{z}_{0}$ is masked. The settings in Eq.17 have been validated to yield high-quality reconstruction results, as shown in Sec.4.4, providing strong evidence of the method’s ability to preserve the features of the initial image.
Furthermore, to minimize the number of sampling steps, improving the accuracy of the estimated $\boldsymbol{\hat{z}}_{0}$ is crucial. We employ the LCM solver Luo et al. (2023a) to predict $\boldsymbol{\hat{z}}_{0}$ from the noised input image $\boldsymbol{z}_{N}$ less than 4 steps, where the LCM is distilled from the models based on the DDIM solver. Experimental results show that the superior denoising capabilities of the LCM solver significantly accelerate the convergence rate and produce more accurate $\boldsymbol{\hat{z}}_{0}$ estimates that align with the target prompt. The measurement characteristics $\boldsymbol{y}$, as defined in Eq.17, involve randomly masking each element of $\boldsymbol{z}_{0}$ with a probability $p$. Since one of the optimization terms focuses on only a small portion of the initial image, both terms in Eq.16 guide the gradient descent in the same direction. As the sampling process progresses, the edited $\boldsymbol{x}_{tgt}$ gradually inherits features from both $\boldsymbol{x}_{0}$ and the target prompt by selectively replacing the necessary attributes. The experimental results of different settings for the optimization defined in Eq.16 are presented in Sec.4.4. The rest of the process mirrors the reconstruction phase, allowing us to progressively achieve the edited $\boldsymbol{x}_{0}$. In summary, the algorithm’s procedure is detailed in Alg. 1, with implementation specifics provided in App.A.4.
Algorithm 1 Posterior Sampling for Image Editing
1:  Require: Diffusion model $\boldsymbol{\epsilon}_{\theta}$, step size $h$, image $\boldsymbol{x}_{0}$, measurement $\boldsymbol{y}$, weight $w$, target prompt $c_{tgt}$ ,$c_{skip}$, $c_{out}$, encoder ${\mathcal{E}}$, decoder $\mathcal{D}$, noise schedule $\alpha(t)$, $\sigma(t)$, multi-step consistency models sequence $\{t_{i}\}_{i=1}^{n}$ and $\{\tau_{i}\}_{i=1}^{N}$.
2:  $\boldsymbol{z}_{0}\sim\mathcal{E}(\boldsymbol{x}_{0})$, $\boldsymbol{z}_{in}=\boldsymbol{z}_{0}$.
3:  Sample $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$, $\boldsymbol{z}_{N}\sim\mathcal{N}(0,\boldsymbol{I})$.
4:  $\boldsymbol{z}_{0}=k_{\text{skip}}(t)\boldsymbol{z}_{N}+k_{\text{out }}(t)\left(\frac{\boldsymbol{z}_{N}-\sigma_{t}\epsilon_{\boldsymbol{\theta}}(\boldsymbol{z}_{N},\boldsymbol{c_{tgt}},t)}{\alpha_{t}}\right)$.
5:  for $i=N-1$ to 00 do
6:     for $j=n-1$ to 00 do
7:        Sample $\boldsymbol{z}_{j}\sim\mathcal{N}\left(\alpha(t_{j})\boldsymbol{z}_{0},\sigma^{2}(t)\boldsymbol{I}\right)$
8:        $\boldsymbol{z}_{0}=c_{\text{skip }}(t)\boldsymbol{z}_{j}+c_{\text{out }}(t)\left(\frac{\boldsymbol{z}_{j}-\sigma_{t}\epsilon_{\boldsymbol{\theta}}(\boldsymbol{z}_{j},\boldsymbol{c_{tgt}},t)}{\alpha_{t}}\right)$
9:     end for
10:     return  $\boldsymbol{z}_{0}$
11:     $\boldsymbol{z}_{0}^{0}=\boldsymbol{z}_{0}$.
12:     for $k=0$ to $m$ do
13:        $\boldsymbol{z}_{0}^{(k+1)}=(1-w)\cdot\boldsymbol{z}_{0}^{(k)}+w\cdot\boldsymbol{z}_{in}-h\cdot\left(\frac{\|z_{0}^{(k)}-z_{0}\|^{2}}{2\sigma_{t}^{2}}+\frac{\|\mathcal{A}\left(\boldsymbol{z}_{0}^{(k)}\right)-\boldsymbol{y}\|^{2}}{2m^{2}}\right)+\sqrt{2h}\epsilon$,   $\epsilon\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$
14:     end for
15:     Sample $\boldsymbol{z}_{\tau_{i-1}}\sim\mathcal{N}(\boldsymbol{z}_{0}^{(N)},\sigma_{{\tau_{i-1}}}\boldsymbol{I})$.
16:  end for
17:  $\boldsymbol{x}_{0}=\mathcal{D}\left(\boldsymbol{z}_{0}\right)$
18:  Return $\boldsymbol{x}_{0}$
@@ 4.1 Experiment Setup
To ensure a fair comparison, all experiments were conducted on the PIE-Bench dataset Ju et al. (2024) using a single A100 GPU to evaluate both image quality and inference efficiency. The PIE-Bench dataset comprises 700 images with 10 types of editing, where each image is paired with a source prompt and a target prompt. In our experiments, the resolution of all test images was set to $512\times 512$. For the reconstruction experiments, we set the initial and target prompts to be identical across all test runs. Additional settings are detailed in App. A.4.
@@ 4.2 Quantitative Comparison
To evaluate the performance of our proposed method, a quantitative comparison is presented in Tab. 1. We compare our method against seven recent image editing approaches: NTIMokady et al. (2023), NPI Miyake et al. (2023), PnP Tumanyan et al. (2023), DI Ju et al. (2024), iCD Starodubcev et al. (2024), DDCM Xu et al. (2024), and TurboEdit Deutch et al. (2024). The comparison is based on three key metrics: background consistency, CLIP Radford et al. (2021) similarity, and efficiency, which are used as comparative benchmarks in our evaluation.
@@
PSNR LPIPS×${}_{\times 10^{2}}$ MSE×${}_{\times 10^{3}}$ SSIM×${}_{\times 10^{2}}$ Whole Edited Time
$\uparrow$ $\downarrow$ $\downarrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\downarrow$
@@
Table 1: Quantitative Comparisons of Editing. Results are evaluated on eight different methods, covering the evaluation of background similarity and editing performance. $\boldsymbol{*}$ means this model benefits from the inference superiority of the SDXL-Turbo model and the speed of our method can be further improved since we adopt LCM-SD1.5 as the baseline model.
The experimental results reflect that PostEdit achieves SOTA performance on editing, which are the “Whole” and “Edited” indicators of the CLIP similarity and the results are significantly better than others. For efficiency, our model is highly efficient with a runtime less than 2 seconds. It is worth noting that our runtime is slightly higher than TurboEdit Deutch et al. (2024), which is mainly due to different baselines. Specifically, TurboEdit employs SDXL-Turbo Sauer et al. (2023) while our framework is based on LCM-SD1.5 Luo et al. (2023a). As shown in App. A.6, SDXL-Turbo Sauer et al. (2023) is almost $2.5$ times faster than LCM-SD1.5 Luo et al. (2023a). We believe the efficiency of our framework can be further improved if we adopt a more efficient baseline like SDXL-Turbo. For the aspect of background preservation, our method possesses the best result on MSE. In the following part, more qualitative results can further demonstrate the superiority of our framework in background preservation.
In conclusion, our method demonstrates superior editing capability and efficiency compared to the selected recent approaches, while also effectively preserving background similarity.
@@ 4.3 Qualitative Comparison of Reconstruction and Editing
To demonstrate the superiority of PostEdit in image reconstruction, we present relevant results in Fig. 3. The experiments indicate that our method exhibits the most stable reconstruction capacity and high quality generation ability compared to NTI Mokady et al. (2023), NPI Miyake et al. (2023) and icd Starodubcev et al. (2024), where there is at least one failure case produced by their method. Specifically, the inversion-free method DDCM Xu et al. (2024) fails to faithfully reconstruct the input images, supporting our claim made in Sec. 1. While other methods yield better results in the given cases, they require significantly longer processing times to achieve competitive outcomes. Therefore, our approach offers the best overall performance when considering inference efficiency, stability in generation, and image quality. More reconstruction results are shown in Fig. 10.
To intuitively compare the editing capability, we present a qualitative comparison of eight methods. As illustrated in Fig. 4, PostEdit most effectively highlights the features present in the target prompt, aligning with the quantitative results shown in Tab. 1. Furthermore, the visualized experiments indicate that our method successfully preserves the initial features. Additional comparison results can be found in App. A.7.
@@ 4.4 Ablation Study
In this section, we conduct various ablation studies and present the results to demonstrate the effectiveness of our framework. (a) We remove the optimization component shown in Eq. 16 and directly apply the adopted SDE/ODE solver to estimate $\boldsymbol{x}_{0}$. The experimental results indicate that the edited images lack background preservation. For instance, in the slanted bicycle example shown in the first row of Fig.5, the staircase on the left side of the original image is transformed into a car in the edited image. (b) We modify the masked probability of our measurement $\boldsymbol{y}$. Notably, there is no discernible difference between the edited images and the input images. (c) We investigate the influence of Proposition 1 on the experimental outcomes, which highlights the effectiveness of the parameter $w$ concerning background similarities.
Figure 5: Ablation Study Relating to Our Settings. We show the results without the optimization process shown in Eq. 16, the measurement $\boldsymbol{y}$ defined as a mask matrix and $\boldsymbol{z}_{in}$ shown in Proposition 1.
@@ 5 Conclusion And Limitation
In this work, we address the errors caused by the unconditional term in Classifier-Free Guidance (CFG) by introducing the theory of posterior sampling to enhance reconstruction quality for image editing. To the best of our knowledge, our work is the first to integrate the theory of posterior sampling into the image editing task based on large text-to-image diffusion models. By minimizing the need for repeated network inference, our method demonstrates fast and accurate performance while effectively preserving background similarity, as evidenced by the results. Ultimately, our approach tackles three key challenges associated with image editing and showcases state-of-the-art (SOTA) performance in terms of editing capabilities and inference speed.
Limitation: Since this method is conditioned on a given text prompt, it can be challenging to represent specific scenes. For example, describing ”a man raising his hand” is considerably more difficult compared to the input format used in ControlNet-related works. Additionally, the quality and speed of generation are contingent upon the baseline models. We are actively considering extending our method to incorporate SDXL-turbo or other state-of-the-art models in the future.
@@ Appendix A Appendix
A.1 Classifier free diffusion guidance

### v2

@@ 3.2 Posterior Sampling for Image Editing
In this section, we present details of the posterior sampling process for high-quality image editing using the DDIM solver Luo et al. (2023a), as outlined in Eq. 6. Unlike the ODE solver used in the image reconstruction task described in Sec. 3.1, image editing requires the solver with higher accuracy to estimate $\boldsymbol{\hat{z}}_{0}$. The measurement $\boldsymbol{y}$ for image reconstruction and editing is defined as
@@
where $\boldsymbol{P}\in\{0,1\}^{n\times p}$ represents a masking matrix composed of elementary unit vectors. This measurement setup not only serves as a specialized configuration for image editing but also demonstrates its capacity to deliver high-quality image reconstruction results, even when $\boldsymbol{z}_{0}$ is masked. The settings in Eq. 17 have been validated to yield high-quality reconstructions, as shown in Sec. 4.3, demonstrating the method’s effectiveness in preserving the features of the initial image.
Furthermore, to minimize the number of sampling steps, improving the accuracy of the estimated $\boldsymbol{\hat{z}}_{0}$ is crucial. According to the experimental results presented in LCM, the superior denoising capabilities of the LCM solver Luo et al. (2023a) are demonstrated to surpass those of the DDIM solver in both speed and accuracy. Consequently, we utilize the LCM solver, distilled from models based on the DDIM solver, to markedly improve the convergence rate and produce more accurate $\boldsymbol{\hat{z}}_{0}$ estimates that closely align with the target prompt in fewer than four steps. The measurement characteristics $\boldsymbol{y}$, as defined in Eq. 17, involve randomly masking each element of $\boldsymbol{z}_{0}$ with a given probability. Since one of the optimization terms focuses on only a small portion of the initial image, both terms in Eq. 16 guide the gradient descent in the same direction. As the sampling process progresses, the edited $\boldsymbol{x}_{tgt}$ gradually inherits features from both $\boldsymbol{x}_{0}$ and the target prompt by selectively replacing the necessary attributes. The experimental results of different settings for the optimization defined in Eq. 16 are presented in Sec. 4.4. The rest of the process mirrors the reconstruction phase, allowing us to progressively achieve the edited $\boldsymbol{x}_{0}$. In summary, the algorithm’s procedure is detailed in Alg. 1, with implementation specifics provided in Appendix A.2.
@@ 4.1 Experiment Setup
To ensure a fair comparison, all experiments were conducted on the PIE-Bench dataset Ju et al. (2024) using the same parameter settings specified in Appendix A.2 and a single A100 GPU to evaluate both image quality and inference efficiency. The PIE-Bench dataset comprises 700 images with 10 types of editing, where each image is paired with a source prompt and a target prompt. In our experiments, the resolution of all test images was set to $512\times 512$. For the reconstruction experiments, we set the initial and target prompts to be identical across all test runs. Additional settings, including forward operators, are provided in the Appendix A.2.
@@ 4.2 Quantitative Comparison
Image Reconstruction.
The methods have special design for image reconstruction are compared: NTI Mokady et al. (2023), NPI Miyake et al. (2023), iCD Starodubcev et al. (2024) and DDCM Xu et al. (2024). The results of quantitative comparison are shown in Tab. 1. Although NTI and NPI achieve better performance on the listed metrics, their computational time costs are substantially higher, exceeding ours by at least an order of magnitude. Compared to the highly efficient inversion-free method, DDCM, PostEdit demonstrates significantly superior performance.
Method Background Preservation Efficiency
PSNR LPIPS×${}^{\downarrow}_{\times 10^{2}}$ MSE×${}_{\times 10^{3}}^{\downarrow}$ SSIM×${}_{\times 10^{2}}^{\uparrow}$ Time
NTI 25.58 7.98 4.37 77.02 $\sim$ 120s
NPI 24.66 9.11 4.73 76.14 $\sim$ 15s
iCD 19.64 17.13 13.50 66.48 $\sim$ 1.8s
DDCM 18.00 17.74 18.94 64.01 $\sim$ 2s
Ours 24.39 9.00 4.75 72.74 $\sim$ 1.5s
Table 1: Quantitative Comparisons of Image Reconstruction. All of the comparison methods include strategies specifically designed for image reconstruction.
@@
PSNR LPIPS×${}^{\downarrow}_{\times 10^{2}}$ MSE×${}_{\times 10^{3}}^{\downarrow}$ SSIM×${}_{\times 10^{2}}^{\uparrow}$ Whole Edited Time
@@
Table 2: Quantitative Comparisons of Image $\boldsymbol{*}$’ indicates models that benefit from SDXL-Turbo’s improved inference.
Image Editing.
We compare our method against recent inversion-based and training-based image editing approaches: NTI, NPI, PnP Tumanyan et al. (2023), DI Ju et al. (2024), iCD, DDCM, TurboEdit Deutch et al. (2024), SPD Li et al. (2024) and GR Titov et al. (2024), IP2P Brooks et al. (2023), OmniGen Xiao et al. (2024) and SeedX Ge et al. (2024) [^1] See Appendix A.12 to get instructions for IP2P, OmniGen, and SeedX based on the input and target prompt.. The comparison is evaluated from three aspects: background consistency, CLIP Radford et al. (2021) similarity, and efficiency. The experimental results shown in Tab. 2 reflect that PostEdit achieves SOTA performance on editing, which are the “Whole” and “Edited” metrics of the CLIP similarity and the results are significantly better than others. For efficiency, our model is highly efficient with a runtime less than 2 seconds. It is worth noting that our runtime is slightly higher than TurboEdit Deutch et al. (2024), which is mainly due to different baselines. Specifically, TurboEdit employs SDXL-Turbo Sauer et al. (2023) while our framework is based on LCM-SD1.5 Luo et al. (2023a). As shown in Appendix A.11, SDXL-Turbo Sauer et al. (2023) is almost $2.5$ times faster than LCM-SD1.5 Luo et al. (2023a). We believe the efficiency of our framework can be further improved if we adopt a more efficient baseline like SDXL-Turbo. In terms of background preservation, our method achieves the best MSE result among all methods with a runtime of less than 2 seconds. The following section presents additional qualitative results, further illustrating the superiority of our framework in editing capabilities and background preservation, while maintaining high efficiency.
@@ 4.3 Qualitative Comparison of Reconstruction and Editing
Image Reconstruction.
We present results of qualitative comparison in Fig. 3. The experiments indicate that PostEdit demonstrates greater robustness and high quality generation ability compared to NTI Mokady et al. (2023), NPI Miyake et al. (2023) and iCD Starodubcev et al. (2024). Specifically, we compared the reconstruction quality across four distinct categories of images: single-object images, complex backgrounds, multi-object scenes, and cartoon images. The inversion-free method DDCM Xu et al. (2024) fails to faithfully reconstruct the input images, supporting our claim made in Sec. 1. While other methods yield better results in the given cases, they require significantly longer processing times to achieve competitive outcomes. Therefore, our approach offers the best overall performance when considering inference efficiency, stability in generation, and image quality. More reconstruction results on complex objects are shown in Fig. 15 and Fig. 16.
Image Editing.
The qualitative comparisons of the image editing results are shown in Fig. 4. The effects of text insertion, deletion, and substitution are provided. PostEdit effectively highlights the features present in the target prompt, which aligns with the quantitative results shown in Tab. 2. To present these findings more clearly, we selected the best-performing classical methods, and their results are shown in Tab. 2. For a comparison of the other baseline methods, please refer to Fig. 12 in Appendix A.7. Additionally, the visualized experiments demonstrate that our method successfully preserves the original features. More comparison results can be found in Appendix A.7.
@@
Figure 5: Ablation Studies. We show the results without the optimization process shown in Eq. 16, the measurement $\boldsymbol{y}$ defined in Eq. 17 and $\boldsymbol{z}_{in}$ shown in Proposition 1.
@@ 4.4 Ablation Study
In this section, we conduct various ablation studies and present the results to demonstrate the effectiveness of our framework. (a) We remove the optimization component shown in Eq. 16 and directly apply the adopted SDE/ODE solver to estimate $\boldsymbol{x}_{0}$. The experimental results indicate that the edited images lack background preservation. For instance, in the slanted bicycle example shown in the first row of Fig. 5, the staircase on the left side of the original image is transformed into a car in the edited image. (b) We modify the masked probability of our measurement $\boldsymbol{y}$. Notably, there is no discernible difference between the edited images and the input images. (c) We investigate the influence of Proposition 1 on the experimental outcomes, which highlights the effectiveness of the parameter $w$ concerning background similarities. Additionally, the quantitative results, as detailed in Tab. 5, highlight the adopted configurations of PostEdit achieve optimal generation performance.
@@ 5 Conclusion And Limitation
In this work, we address the errors caused by the unconditional term in Classifier-Free Guidance by introducing the theory of posterior sampling to enhance reconstruction quality for image editing. To the best of our knowledge, our work is the first to integrate the theory of posterior sampling into the image editing task based on large text-to-image diffusion models. By minimizing the need for repeated network inference, our method demonstrates fast and accurate performance while effectively preserving background similarity, as evidenced by the results. Ultimately, our approach tackles three key challenges associated with image editing and showcases state-of-the-art performance in terms of editing capabilities and inference speed.
Limitation: PostEdit faces challenges in representing highly specific scenes. For example, describing ”a man raising his hand” is considerably more difficult compared to the structured input formats used in ControlNet-related methods. Furthermore, its ability to maintain background consistency is limited and requires improvement. Additionally, the quality and speed of generation are strongly influenced by the performance of the underlying baseline models.
@@ Appendix A Appendix
A.1 Related Work in Image Editing
Previous works in image editing can be broadly categorized into two paradigms: inversion-based and training-based.
Inversion-based Methods.
Several works Zhang et al. (2023d); Ruiz et al. (2023); Gal et al. (2023) focus on modifying the training process of diffusion models to incorporate the information from the initial images. Specifically, the optimization process can be operated in two different spaces: textual space and model space. In textual space, a set of methodologies Dong et al. (2022); Valevski et al. (2023); Han et al. (2023) aim to optimize textual embeddings to perform various editing tasks effectively. In model space, a research line Ruiz et al. (2023); Qiao et al. (2024); Avrahami et al. (2023) intends to further updates modules from the base model to enhance reconstruction capabilities. For example, several studies Kawar et al. (2023); Shi et al. (2024); Zhang et al. (2023c) optimize both textual embeddings and model parameters to ensure content consistency following non-rigid editing or localized distortions. For forward-based inversion, it can also be divided into two categories, which is DDIM inversion and DDPM inversion. Previous works like Mokady et al. (2023) and Miyake et al. (2023) are designed to approximate the inversion trajectory to deal with accumulated errors. Dong et al. (2023) focuses on optimizing the text embedding, which is then interpolated with the target embedding during the editing process. Some approaches Han et al. (2024); Cho et al. (2024) aim to bypass the time-consuming optimization processes of the aforementioned methods while preserving their reconstruction capabilities. Inspired by normalizing flow models Dinh et al. (2015; 2017), EDICT Wallace et al. (2023) reformulates DDIM processes by simultaneously tracking two associated noisy variables at each step during inversion. These variables can be exactly derived from one another during the sampling phase.
Training-based Methods.
Since some advanced approaches in zero-shot or few-shot settings require time-consuming optimization Ruiz et al. (2023); Gal et al. (2023); Mokady et al. (2023) or are highly sensitive to hyperparameters Cao et al. (2023); Kawar et al. (2023); Hertz et al. (2023); Huberman-Spiegelglas et al. (2024), several studies Brooks et al. (2023); Ye et al. (2023) aim to train task-specific models with substantial amounts of data to directly transform source images into target images under user guidance. Instruction-based editing Brooks et al. (2023); Zhang et al. (2023a); Xie et al. (2023) provides an intuitive approach for image manipulation, allowing users to input command-style text instead of providing an exhaustive description. For image inpainting, a group of methods Huang et al. (2024); Wang et al. (2023) focuses on completing missing parts of an image under text guidance. Additionally, image translation Isola et al. (2017); Zhang et al. (2023b) seeks to transform the source image into a target domain, such as converting night to daytime or sketch to a natural image. Another type of training-based method is content-free editing, which Ruiz et al. (2023); Wei et al. (2023) aims to preserve the high-level semantics of the source images in the final results. Content-free editing can be further categorized into subject-driven customization and attribute-driven customization. Subject-driven customization Wei et al. (2023); Li et al. (2023); Arar et al. (2023); Chen et al. (2023) is designed to capture the identity of the target and generate novel images that place it in new contexts. In contrast, attribute-driven customization Lee et al. (2024) focuses on extracting and manipulating attributes in a more fine-grained manner.
A.2 Implementation Details
The main hyper-parameters of the PostEdit are briefly summarized in Tab. 3.
SD Model.
We adopt LCM-SD1.5 for all the experiments Luo et al. (2023a).
Parameters of Consistency models.
$c_{skip}$ and $c_{out}$ shown in line 6 of Alg. 1 are set to 0 and 1 for most cases respectively.
Hyper-parameters in Alg. 1.
$N$ is set to $5$ for schedule $\{\tau_{i}\}_{i=0}^{N-1}$. To ensure higher efficiency and quality at the same time, $\boldsymbol{z}_{N}$ is sampled through
where $t$ is set to $501$ generally following the DDPM scheduler Ho et al. (2020). Additionally, $n$ is set to 1 for the sequence of diffusion sampler $\{t_{j}\}_{j=0}^{n-1}$ to further improve inference speed. The parameter $T$ shown in line 16 of Alg. 1 is set to 100 to ensure optimal quality. For Eq. 16, $m$ is set as $0.01$ for both the reconstruction and editing task while $\sigma_{t}$ corresponds to the timestep of the DDPM scheduler. Generally, we apply Eq. 8 for $1$ step to estimate $\boldsymbol{z}_{0}$, and then according to the following schedule to make $\boldsymbol{z}_{\tau_{i}}$ progressively converge to $\boldsymbol{z}_{0}$.
$\{\tau_{i}\}_{i=1}^{5}=\{501,401,301,201,101,1\}.$ (19)
The parameter $w$ is usually set to a minimal value such as $0.1$ for most cases or 00 and $0.2$ for easy and hard cases.Additionally, $h$ is initially set to 1e-5 for image editing and reconstruction tasks. It is dynamically adjusted at each recursion step, as described in lines 9 to 12 of Alg. 1, using the following equation:
$h=\left(1+\frac{k}{T}\cdot(0.01-1)\right)\cdot h,$ (20)
where $k$ and $T$ are the same with definition of line 9 of Alg. 1.
ODE Solvers.
We adopt the solver of LCM Luo et al. (2023a) distilled from Dreamshaper v7 fine-tune of Stable-Diffusion v1-5 for images editing task. For reconstruction, different solvers, for instance, DDIM Song et al. (2021a), DDPM Ho et al. (2020), and Song et al. (2023) based on Stable Diffusion Rombach et al. (2022) are able to work out satisfied reconstruction quality.
Probability of Masked Features.
We use the probability equal to 0.5 for a randomly mask process, which represents whether one of the latent features is masked or not.
Measurement $\boldsymbol{y}$ Used for Better Quality of Image Reconstruction.
The measurement $\boldsymbol{y}$ can be chosen from the following Eq. 21 to further improve the reconstruction ability of PostEdit, which are defined as linear and nonlinear operations relating to initial image $\boldsymbol{z}_{0}$ in latent space
where $\boldsymbol{F}$ and $\boldsymbol{P}$ denote the 2D discrete Fourier transform matrix and the oversampling matrix with ratio $k/n$ respectively for Eq. 21. However, the forward operator term shown in Eq. 21 reflects poor editing ability, and all our editing and reconstruction results all based on the measurement shown in Eq. 17.
Oversampling Matrix.
We set $\sigma$ shown in Eq. 21 to $0.01$ and use an oversampling factor $k=2$ and $n=8$.
2D Discrete Fourier Transform Matrix.
The 2D Fourier transform is defined as
$u=0,1,...,M-1;\ \ v=0,1,...,N-1,$
FFHQ Model.
We adopt the ffhq_10m.pt with a size of 357.1MB as the baseline model for all the experiments relating to the FFHQ dataset.
A.3 Hyperparameter Sensitivity Analysis
Notation Values Description
$f$ 0.5 Appearance probability of 00 in Matrix $P$ shown in Eq. 17
Optimization Steps 100 Operating steps of Eq. 16
$w$ 0.1 The weighting coefficient of Proposition. 1
Table 3: Main Hyper-parameters.
Method Background Preservation CLIP Similarity Efficiency
PSNR LPIPS×${}^{\downarrow}_{\times 10^{2}}$ MSE×${}_{\times 10^{3}}^{\downarrow}$ SSIM×${}_{\times 10^{2}}^{\uparrow}$ Whole Edited Time
$f=0.3$ 27.20 6.09 2.91 82.77 25.93 22.40 $\sim$1.5s
$f=0.7$ 24.43 12.16 6.06 77.64 26.73 24.28 $\sim$1.5s
Optimization Steps $=50$ 25.49 9.39 4.85 79.95 26.61 23.59 $\sim$1.1s
Optimization Steps $=150$ 26.59 7.19 3.77 82.05 26.51 23.47 $\sim$1.8s
$w=0.3$ 27.00 8.50 4.39 82.70 26.34 23.49 $\sim$1.5s
$w=0.5$ 27.75 5.47 2.84 83.61 25.83 22.19 $\sim$1.5s
$w=0$ 24.01 10.92 6.24 77.28 26.45 23.45 $\sim$1.5s
Ours Default 27.04 6.38 3.24 82.20 26.76 24.14 $\sim$1.5s
Table 4: Quantitative Results of Hyperparameter Sensitivity Analysis
The hyperparameters used for PostEdit are listed in Tab.3, and its performance under different settings is presented in Tab.4. We conduct the following hyperparameters sensitivity analysis:
  • Appearance Probability of 0 in Matrix $\boldsymbol{P}$. A higher probability (e.g., 0.7) improves CLIP Similarity but degrades Background Preservation metrics (e.g., PSNR and SSIM). Conversely, a lower probability (e.g., 0.3) favors background preservation at the expense of CLIP similarity.
  • Optimization Steps. Reducing the number of steps, such as 50, decreases computation time but negatively impacts performance across most metrics. Increasing the steps to 150 offers slight performance improvement but reduces efficiency. The chosen configuration of 100 steps strikes a balance between quality and runtime.
  • Weighting Coefficient. Setting $w$ to zero results in poor performance for both background preservation and editing capabilities. While increasing $w$ enhances background consistency, editing performance remains suboptimal.
  • Our Configuration. The default settings strike a balance across all metrics, achieving competitive results in background preservation, editing alignment, and efficiency.
A.4 Comparison The Images Layout of Different Datasets
A.5 Quantitative Results of Ablation Study
To better reflect the effectiveness of our adopted settings, we also conduct a quantitative results of the ablation study shown in Tab. 5. The results further verify the performance of all settings of PostEdit.
Method Background Preservation CLIP Similarity Efficiency
PSNR LPIPS×${}^{\downarrow}_{\times 10^{2}}$ MSE×${}_{\times 10^{3}}^{\downarrow}$ SSIM×${}_{\times 10^{2}}^{\uparrow}$ Whole Edited Time
No Posterior Sampling 21.31 16.88 10.36 73.21 26.38 23.28 $\sim$1s
No mask 28.31 4.61 2.64 84.15 25.17 20.94 $\sim$1.5s
No $\boldsymbol{z_{in}}$ 24.01 10.92 6.24 77.28 26.45 23.45 $\sim$1.5s
Ours Full 27.04 6.38 3.24 82.20 26.76 24.14 $\sim$1.5s
Table 5: Quantitative Comparisons of Ablation Study.
A.6 Results of Long-text Editing
We conducted long-text editing experiments, with the results presented in Fig. 8, Fig. 9, Fig. 10 and Fig. 11. These results demonstrate that the editing capabilities of PostEdit extend beyond simple word replacements.
A.7 More Editing Results
Here, we qualitatively compare with other baselines including NPI, GR, DI, and IP2P, SeedX. The results are shown in Fig. 12. The results support the conclusion in the manuscript.
A.8 Reconstruction Results of Different Forward Operators
Images in Fig. 14 are reconstructed through different forward operators as shown in Eq. 17 and Eq. 21. The corresponding quantitative comparison of image reconstruction of different measurements is shown in Tab. 6. The corresponding quantitative comparison of image reconstruction of different measurements is shown in Tab. 6.
Measurement Background Preservation Efficiency
PSNR LPIPS×${}^{\downarrow}_{\times 10^{2}}$ MSE×${}_{\times 10^{3}}^{\downarrow}$ SSIM×${}_{\times 10^{2}}^{\uparrow}$ Time
Eq. 21 24.90 7.60 4.31 74.03 $\sim$ 1.5s
Eq. 17 (Used) 24.39 9.00 4.75 72.74 $\sim$ 1.5s
Table 6: Quantitative Comparisons of Image Reconstruction using different Measurements.
A.9 Intermediate Results
The intermediate state for different iterative steps are shown detailed in Fig. 13.
A.10 Additional Reconstruction Results
The additional reconstruction results are exhibited in Fig. 15 and Fig. 16. Additionally, We compare reconstruction quality of different methods show in Fig. 19. The results reflect the effectiveness of PostEdit to reconstruct high frequency information.
A.11 Comparison Between LCM-SD1.5 And SDXL-Turbo
A.12 Generating edit instruction for comparing with instruction-based image editing approaches
We provide GPT4-o with the following instruction shown in Fig. 18 to convert the differences between the input prompt and the edited prompt into editing instructions suitable for IP2P and SeedX.
A.13 User Study
We invited 34 anonymous volunteers to rank the preferred results of image editing results. The results are evaluated by the quality of background preservation and features aligned with the given target prompt. The feedback is shown in Tab. 7 and Tab. 8 and the preference represent a vote of the participants. The results indicate that PostEdit outperforms the compared baselines and is the most popular approach for both image reconstruction and editing tasks.
Method Ours iCD DI SPD NTI PnP DDCM
Preference (Editing) 81 22 8 23 5 13 12
Method TurboEdit NPI GR IP2P SeedX OmniGen
Preference (Editing) 24 18 30 1 1 55
Table 7: User Study of Image Editing.
Method Ours NTI NPI DDCM iCD
Preference (Reconstruction) 101 66 56 1 11
Table 8: User Study of Image Reconstruction.
A.14 Classifier free diffusion guidance